# DeltaZip: Efficient Serving of Multiple Full-Model-Tuned LLMs

### v2

@@ 1. Introduction
While fine-tuning is typically a one-time effort performed off the critical path, LLM serving is critical to optimize as it is typically recurring and latency-critical. Techniques such as continuous batching (yu_orca_2022, ), paged attention (kwon_efficient_2023, ), prompt processing disaggregation (strati_dejavu_2024, ; patel_splitwise_2024, ; zhong_distserve_2024, ), and tensor parallelism (li_alpaserve_2023, ) optimize inference latency and throughput for individual models, however concurrently serving many model variants presents additional challenges.
Figure 1 shows that the requests to each model variant in an LLM-based chatbot service are sporadic, making it difficult to batch requests and decide how to provision GPU resources for each model. Dedicating GPUs for each model variant minimizes latency, but requires many expensive GPUs that would sit idle for long periods of time (yellow areas in Figure 1). On the other hand, swapping model variants on a limited pool of GPUs reduces cost and improves utilization, but adds latency on the critical path of requests. Optimizing the storage hierarchy to reduce swapping latency can help (fu2024serverlessllm, ), but requests still experience long queuing delays due to limited batching opportunities when treating each model variant as a separate model.
The state-of-the-art approach to LLM serving with the proliferation of fine-tuned model variants is to adopt an entirely new fine-tuning paradigm, parameter-efficient tuning (PEFT). Instead of tuning model weights directly, PEFT tunes a compact, adjunctive model adapter (i.e., a small set of extra parameters) that is attached to the model for serving. For example, low-rank adaptation (LoRA) (hu_lora_2021, ) is a popular PEFT method that freezes models weights and attaches low-rank matrices to the model structure which are fine-tuned on task-specific data. Systems like Punica (chen2024punica, ) and S-LoRA (sheng_s-lora_2023, ) leverage the small size of LoRA adapters and the common base model weights to efficiently swap adapters and batch requests to optimize inference latency and throughput.
@@ 3.1. Opportunities and Key Insights
3. GPU support for delta computations. We can accelerate delta inference by leveraging features of modern GPU hardware, such as: 1) using sparse tensor cores (bai_structured_2023, ) for sparse delta computations, 3) performing multiple low-precision matrix multiplications in parallel to improve stream multi-processors (SMs) utilization, and 3) reducing memory bandwidth pressure for delta matrix computations by designing a custom GPU kernel that minimizes data movement from GPU global memory to device memory.
@@ 3.2. System Overview
To realize the above opportunities in practice, we propose DeltaZip. Our key contributions are a model delta compression algorithm ($\Delta$Compress) and the end-to-end DeltaZip serving system. The algorithm compresses FMT deltas into a hardware-efficient, low-precision, and sparse format that preserves high accuracy, while the serving system incorporates state-of-the-art serving optimizations and adapts them for low-latency, high-throughput inference with model deltas.
@@ 4. Model Delta Compression
When a model developer registers an FMT model with DeltaZip, the Delta Compressor computes the delta and applies a compression pipeline (§4.1) to reduce the size of the model state. The compression pipeline only runs when a developer registers a new model. We develop the $\Delta$Compress algorithm (§4.2) to apply quantization and sparsification in the compression pipeline in a way that preserves high accuracy.
@@ 4.1. DeltaZip Compression Pipeline
Step 2 applies structured 2:4 pruning (zhou_learning_2021, ; hubara_accelerated_2021, ), which involves setting at least two elements among each group of four contiguous elements to 0 in the delta matrix. After structured pruning, DeltaZip only needs to store the non-zero values of the delta matrix and their 2-bit indices. Compared to applying quantization only or unstructured sparsity, structured sparsity enables higher peak performance with large input size. Figure 6 shows a microbenchmark of matrix multiplication performance with different compression configurations. We observe that with small input sizes (e.g., from 1 to 4, which is common during the decode phase), structured sparse matrix multiplication achieves similar performance to quantization-only compression and outperforms the uncompressed version. This is mainly because both sparsity and quantization reduce data movement between GPU HBM and compute unit. However, with large input sizes (e.g., from 16 to 4096, which is typical during the prefill phase), structured sparsity significantly outperforms quantization-only compression. This is because structured sparsity leverages the sparse tensor cores on GPUs to achieve higher performance.
@@ 4.2. $\Delta$Compress Algorithm
The major distinction from full model compression is that $\Delta$Compress reconstructs the weight matrix for each layer (Line 6) after compressing the delta and computes the input for next layer (Line 7). This is because the input data (i.e., $\mathbf{X_{n}}$ in Line 5) is crucial for the compression algorithm. Without re-adding the base weight matrix and reconstructing the weight matrix for each layer, the magnitude of the deltas causes diminishing outputs, leading to vanishing activations in deeper layers. The vanishing activations, as the input to the next layer, will make the compression algorithm fail to capture the input (i.e., $X_{n}$ Line 5). By extracting the delta and reconstruting the weight for each layer on the fly, $\Delta$Compress ensures proper calibration and maintains high model quality.
Beyond model quality, $\Delta$Compress has two advantages: 1) Low memory requirement. As $\Delta$Compress performs layer-wise compression, it only needs a GPU with sufficient memory for a single layer to perform the forward pass. 2) No need to retrain the model. Unlike some other compression algorithms (chai_int21_2023, ; tseng2024quip, ), $\Delta$Compress does not require further fine-tuning the model to recover the model quality.
@@ 5.1. Base and Delta Decoupling
DeltaZip’s serving engine always keeps the base model in GPU memory, [^1] If there are $M$ base models, we divide the GPU cluster into $M$ sets of GPUs, each dedicated to serving a particular base model’s variants. and swaps compressed deltas on-demand to serve inference requests. The naive approach to serve a fine-tuned model is to load its compressed delta, decompress it, add it to the base model, and then perform inference. However, this approach is inefficient for several reasons: 1) it requires decompressing the delta on the critical path, which adds latency; 2) it does not allow batching requests to different model variants with the same base model; 3) performing inference after adding the delta back to the full-precision base model does not leverage low-precision computation to reduce delta inference latency; and 4) storing decompressed deltas in GPU memory to add back to the base model limits how many deltas can fit concurrently in the GPU.
@@ 5.4. Continuous Batching and Scheduling with Delta
Last but not least, DeltaZip optimizes inference by extending continuous batching (yu_orca_2022, ), a standard technique to improve GPU utilization for LLM serving. DeltaZip implements a software scheduler per set of GPU workers that form a tensor parallel serving group. At every iteration, the scheduler picks $N$ models to serve concurrently on a first-come-first-served basis, where $N$ is at most the number of deltas that can fit concurrently in GPU memory across the workers. After selecting the first $N$ deltas, the scheduler searches the request queue for all requests that belong to the selected $N$ models. To maximize batching, it allows requests to skip the line if they are for one of the $N$ selected models at the head of the queue. The scheduler delivers the batch of requests to the Model Runner, which starts to load the requested deltas, performs prompt processing if needed, and then batches request decoding for different models.
Tuning $N$, the number of concurrent deltas. DeltaZip tunes $N$ empirically based on a short trace profiling phase. We explore the impact of varying $N$. Intuitively, processing more deltas concurrently (larger $N$) allows for more extensive request batching, enhancing throughput. However, collocating many deltas increases GPU memory pressure, potentially degrading performance if there are already many requests to batch for a particular delta. On the left side of Figure 10, we show how $N$ affects the serving latency when executing a 25 second time interval of a trace with arrival rate $3$ and zipf-4.0 model popularity distribution. In this case, DeltaZip would pick $N=3$ as it achieves optimal performance during profiling. We then run a series of traces under different settings. The right side of Figure 10 shows that $N=3$ remains optimal or near-optimal across different arrival rates and model popularity distributions. We generally find that offline profiling with a short trace is sufficient to determine the (near-)optimal number of concurrent deltas. Dynamic tuning can also be implemented.
@@ 6.1. Performance Evaluation Setup
Workload traces. To evaluate the serving performance, we use the prompts and responses sampled from the LMSys Chatbot Arena (zheng_judging_2023, ). We consider three types of model popularity distribution: 1) Uniform: all models are equally popular. 2) Skewed: model popularity follows a Zipf-$\alpha$ distribution. In other words, the popularity of the $i$-th model is proportional to $1/i^{\alpha}$. We set $\alpha=1.5$ for the skewed distribution. 3) Azure trace: since there is no publicly available traces for multi-variant LLM serving available, following previous work (li_alpaserve_2023, ), we use the Azure serverless function traces (shahrad_serverless_2020, ; zhang_faster_2021, ) as a proxy. Note the traffic in this trace is highly bursty and the model distribution is highly skewed. Except for the small-scale timeline analysis in Figure 16 (which we show for illustration purposes) and ablation studies, we run the traces for 5 minutes under different arrival rates and model distribution. Unless otherwise stated, we assume there are 32 model variants that need to be served.
Metrics. We use downstream accuracy to evaluate the post-compression quality and three metrics for serving performance: 1) average end-to-end (E2E) latency, 2) average time to first token (TTFT), which is an important indicator of the system’s responsiveness, 3) throughput and 4) SLO attainment (i.e., percentage of requests served within a given SLO requirement) for TTFT and E2E Latency.
@@
Table 1. Model quality of DeltaZip vs. uncompressed (FP16) vs. SparseGPT (frantar_sparsegpt_2023, ) and AWQ (AWQ, ). For Pythia, T1, T2, T3 show the accuracy on 3 downstream tasks (Amazon Review Classification, Synthetic Palindrome Numbers, Yes/No Question) from natural instructions (mishra_cross-task_2022, ). For other models, T1, T2, T3 show accuracy on 3 standard benchmarks (BoolQA, TruthfulQA, LogiQA) in lm-eval-harness (eval-harness, ). $\star$ indicates 50% structured pruning and quantization.

### v3

@@ Abstract.
Machine Learning Systems, Model Compression, Quantization, Sparsity, Hardware Acceleration, Model Inference and Serving
copyright: rightsretainedjournalyear: 2025conference: Twentieth European Conference on Computer Systems; March 30–April 3, 2025; Rotterdam, Netherlandsbooktitle: Twentieth European Conference on Computer Systems (EuroSys ’25), March 30–April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3689031.3717468isbn: 979-8-4007-1196-1/25/03ccs: Computer systems organization Cloud computingccs: Computing methodologies Machine learning
@@ 1. Introduction
While fine-tuning is typically a one-time effort performed off the critical path, LLM serving is critical to optimize as it is typically recurring and latency-critical. Techniques such as continuous batching (yu_orca_2022, ), paged attention (kwon_efficient_2023, ), prompt processing disaggregation (strati_dejavu_2024, ; patel_splitwise_2024, ; zhong_distserve_2024, ), and tensor parallelism (li_alpaserve_2023, ) optimize inference latency and throughput for individual models. However, the growing need to serve many model variants concurrently presents additional challenges. The need is particularly pronounced for LLM service providers, such as DeepInfra (deepinfracustommodels, ), Together.AI (togetheraicustommodels, ), or Fireworks.AI (fireworksondemand, ), who serve not only the popular base model, but also various fine-tuned models from either customers or the community.
On these service providers, the popularity of each model varies significantly between base models and their fine-tuned variants. From OpenRouter statistics (openrouter_stats, ), while the base model (Llama-3.1-70B) consumes $\sim$1 billion tokens per day, the deployed fine-tuned variants consume much less, e.g., $\sim$100 million for Nemotron-70B and Hermes-3-70B, and less than 10 million for Sonar-70B and Lumimaid-70B. In addition, Figure 1 shows that requests to base models and their variants can range from sporadic to very dense. For example, the models alpaca-13b (alpaca_model, ), koala-13b (koala_model, ), gpt4all-13b (gpt4all_model, ), vicuna-13b (vicuna_model, ) and wizardlm-13b (wizardlm_model, ) are all fine-tuned variants of the base model llama-13b. We observe how certain variants, such as wizardlm-13b and gpt4all-13b, are highly popular over time while others, such as alpaca-13b and vicuna-13b, have sporadic requests. The varying popularity may be due to different accuracy on certain downstream tasks, and the irregular request patterns make it difficult to batch requests and decide how to provision GPU resources for each model.
Dedicating GPUs for each model variant minimizes latency, but requires many expensive GPUs that would sit idle for long periods of time (yellow areas in Figure 1). On the other hand, swapping model variants on a limited pool of GPUs reduces cost and improves utilization, but adds latency on the critical path of requests. Optimizing the storage hierarchy to reduce swapping latency can help (fu2024serverlessllm, ), but requests still experience long queuing delays due to limited batching opportunities when treating each model variant as a separate model.
For LLM service providers, serving this long tail of less popular models cost-efficiently while maintaining low latency and high quality is a key challenge. The state-of-the-art approach to LLM serving with the proliferation of fine-tuned model variants is to adopt an entirely new fine-tuning paradigm, parameter-efficient tuning (PEFT). Instead of tuning model weights directly, PEFT tunes a compact, adjunctive model adapter (i.e., a small set of extra parameters) that is attached to the model for serving. For example, low-rank adaptation (LoRA) (hu_lora_2021, ) is a popular PEFT method that freezes models weights and attaches low-rank matrices to the model structure which are fine-tuned on task-specific data. Systems like Punica (chen2024punica, ) and S-LoRA (sheng_s-lora_2023, ) leverage the small size of LoRA adapters and the common base model weights to efficiently swap adapters and batch requests to optimize inference latency and throughput. Given the needs for deploying fine-tuned LLMs, these solutions have been integrated into commercial platforms such as LoRAX (predibase_lorax, ).
@@
Our compression pipeline and serving system is open source and available at https://github.com/eth-easl/deltazip.
@@ 3.1. Opportunities and Key Insights
3. GPU support for delta computations. We can accelerate delta inference by leveraging features of modern GPU hardware, such as: 1) using sparse tensor cores (bai_structured_2023, ) for sparse delta computations, 2) performing multiple low-precision matrix multiplications in parallel to improve stream multi-processors (SMs) utilization, and 3) reducing memory bandwidth pressure for delta matrix computations by designing a custom GPU kernel that minimizes data movement from GPU global memory to device memory.
@@ 3.2. System Overview
To realize the above opportunities in practice, we propose DeltaZip. Our key contributions are a model delta compression algorithm ($\Delta$Compress) and the end-to-end DeltaZip serving system. The algorithm compresses FMT deltas into a hardware-efficient, low-precision, and sparse format that preserves high accuracy, while the serving system incorporates state-of-the-art serving optimizations and adapts them for low-latency, high-throughput inference with model deltas. DeltaZip is designed for LLM service providers to operate, and LLM users can interact with the system through the standard HTTP API frontend, as if using a dedicated LLM.
@@ 4. Model Delta Compression
When model developers register an FMT model into the system, the Delta Compressor computes the delta and applies a compression pipeline (§4.1) to reduce the size of the model state. The compression pipeline only runs when a developer registers a new model. We develop the $\Delta$Compress algorithm (§4.2) to apply quantization and sparsification in the compression pipeline in a way that preserves high accuracy.
@@ 4.1. DeltaZip Compression Pipeline
Step 2 applies structured 2:4 pruning (zhou_learning_2021, ; hubara_accelerated_2021, ), which involves setting at least two elements among each group of four contiguous elements to 0 in the delta matrix. With structured pruning, DeltaZip only stores the non-zero values of the delta matrix and their 2-bit indices. The main motivation for applying structured sparsity is due to its hardware-efficiency. Compared to quantization only or unstructured sparsity, the particular pattern we used in structured sparsity enables higher peak performance with large input size, since modern NVIDIA GPUs from the Ampere architecture onwards and AMD RDNA/CDNA GPUs have specialized hardware support for this feature (bai_structured_2023, ; amdsparsitysupport, ). This hardware support is automatically enabled when running compatible kernels. Given the trend toward hardware optimization, we expect future GPUs will continue to support this feature.
Figure 6 shows a microbenchmark of matrix multiplication performance with different compression configurations. We observe that with small input sizes (e.g., from 1 to 4, which is common during the decode phase), structured sparse matrix multiplication achieves similar performance to quantization-only compression and outperforms the uncompressed version. This is mainly because both sparsity and quantization reduce data movement between GPU HBM and compute unit. However, with large input sizes (e.g., from 16 to 4096, which is typical during the prefill phase), structured sparsity significantly outperforms quantization-only compression, as structured sparsity leverages the sparse tensor cores (bai_structured_2023, ) on GPUs to achieve higher performance.
@@ 4.2. $\Delta$Compress Algorithm
The major distinction from full model compression is that $\Delta$Compress reconstructs the weight matrix for each layer (Line 6) after compressing the delta and computes the input for next layer (Line 7). This is because the input data (i.e., $\mathbf{X_{n}}$ in Line 5) is crucial for the compression algorithm. Without re-adding the base weight matrix and reconstructing the weight matrix for each layer, the magnitude of the deltas causes diminishing outputs, leading to vanishing activations in deeper layers. The vanishing activations, as the input to the next layer, will make the compression algorithm fail to capture the input (i.e., $X_{n}$ Line 5). By extracting the delta and reconstructing the weight for each layer on the fly, $\Delta$Compress ensures proper calibration and maintains high model quality.
Beyond model quality, $\Delta$Compress has three advantages: 1) Low memory requirement. As $\Delta$Compress performs layer-wise compression, it only needs a GPU with sufficient memory for a single layer to perform the forward pass.
2) No need to retrain the model. Unlike some other compression algorithms (chai_int21_2023, ; tseng2024quip, ), $\Delta$Compress does not require further fine-tuning the model to recover the model quality. 3) Data-Efficiency. $\Delta$Compress requires a small calibration set to achieve high accuracy, which does not need to be the exact, entire training set. In our experiments, we found $256$ samples chosen from open-source instruction tuning dataset UltraChat (ding2023enhancingchatlanguagemodels, ) are sufficient for calibration. In practice, if the users fine-tune their models through the service provider, then the training data can be directly used as the calibration set. If the users fine-tune their models on their own, they can use a small, representative subset (e.g., by uniform sampling) of the training data as the calibration set. The users can also evaluate the model quality after compression using an evaluation set to ensure the quality meets their requirements, and adjust the calibration set if needed.
Depending on the model size and hardware, the compression process takes some time to run. From our experiments, compressing a 7B model takes around 30 minutes on a single RTX 3090 GPU and AMD EPYC 7313 16-Core CPU. Although this still introduces extra cost and overhead, the process is offline (i.e., not on the critical inference path) and only runs once when the model is registered, making it acceptable.
@@ 5.1. Base and Delta Decoupling
DeltaZip’s serving engine always keeps the base model in GPU memory and swaps compressed deltas on-demand to serve inference requests. The naive approach to serve a fine-tuned model is to load its compressed delta, decompress it, add it to the base model, and then perform inference. However, this approach is inefficient for several reasons: 1) it requires decompressing the delta on the critical path, which adds latency; 2) it does not allow batching requests to different model variants with the same base model; 3) performing inference after adding the delta back to the full-precision base model does not leverage low-precision computation to reduce delta inference latency; and 4) storing decompressed deltas in GPU memory to add back to the base model limits how many deltas can fit concurrently in the GPU.
@@
If there are $M$ base models and $M>1$, we divide the GPU cluster into $M$ sets of GPUs, each dedicated to serving a particular base model and its fine-tuned variants. Hence, even with many model variants, we only need $M$ sets of GPUs to serve all these models. LoRA serving systems (chen_punica_2023, ; sheng_s-lora_2023, ) have a similar assumption. However, operators can also deploy multiple base models per GPU if there is sufficient memory.
@@ 5.4. Continuous Batching and Scheduling with Delta
Last but not least, DeltaZip optimizes inference by extending continuous batching (yu_orca_2022, ), a standard technique to improve GPU utilization for LLM serving. DeltaZip implements a software scheduler per set of GPU workers that form a tensor parallel serving group. At every iteration, the scheduler picks $K$ requests to serve concurrently on a first-come-first-served basis. These $K$ requests belong to at most $N$ deltas that can fit concurrently in GPU memory. After selecting the $K$ requests and $N$ deltas, the scheduler searches the request queue for all requests that belong to the selected $N$ deltas. The intuition behind this design is to maximize batching by allowing requests to skip the line if and only if they are for one of the $N$ selected deltas. The scheduler then delivers the batch of requests to the Model Runner, which starts to load the requested deltas, performs prompt processing if needed, and then batches request decoding for different models.
Tuning $N$, the number of concurrent deltas. DeltaZip tunes $N$ empirically based on a short trace profiling phase. We explore the impact of varying $N$. Intuitively, processing more deltas concurrently (larger $N$) allows for more extensive request batching, enhancing throughput. However, collocating many deltas increases GPU memory pressure, potentially degrading performance if there are already many requests to batch for a particular delta. We conduct a microbenchmark on RTX 3090 GPU and synthetic trace to study the effectiveness of offline profiling. On the left side of Figure 10, we show how $N$ affects the serving latency when executing a 25 second time interval of a trace with arrival rate $3$ and zipf-4.0 model popularity distribution. In this case, DeltaZip would pick $N=3$ as it achieves optimal performance during profiling. We then run a series of traces under different settings. The right side of Figure 10 shows that $N=3$ remains optimal or near-optimal across different arrival rates and model popularity distributions. We generally find that offline profiling with a short trace is sufficient to determine the (near-)optimal number of concurrent deltas. Dynamic tuning can also be implemented. If offline profiling is not possible, operators can heuristically set $N$. If there are only few requests per delta, we can allow more deltas (and hence more requests in parallel) to improve the GPU utilization. If there are many requests per delta, we can reduce $N$ to avoid memory pressure. Like other systems, operators can also adjust the maximum batch size and concurrent deltas as hyperparameters.
@@
Scalability. DeltaZip employs a hierarchical management strategy when the number of deltas exceeds both GPU and host memory capacity. When there are more deltas than the capacity of GPU memory and even host memory, DeltaZip will swap deltas in and out of GPU/CPU memory to disk, and only load deltas when they are needed. For an inference batch, only a fixed and pre-set number of deltas (or less) will be processed concurrently and thus a single batch will not exceed the GPU memory capacity. The performance of DeltaZip will gracefully degrade if there are sporadic requests for deltas that reside on disk due to swapping.
@@ 6.1. Performance Evaluation Setup
FP16 83.70 45.10 27.65 1.00$\times$
SparseGPT (4$\star$) 74.86 38.79 21.50 1.53$\times$
AWQ (4bit) 83.14 44.97 28.11 2.33$\times$
DeltaZip (4$\star$) 83.79 44.24 27.50 3.26$\times$
Gemma 2 2B DeltaZip (2$\star$) 83.66 44.40 26.26 4.08$\times$
FP16 88.87 51.75 33.18 1.00$\times$
SparseGPT (4$\star$) 86.26 45.71 29.95 2.43$\times$
AWQ (4bit) 88.47 51.24 33.33 3.10$\times$
DeltaZip (4$\star$) 88.96 51.32 34.87 4.61$\times$
Gemma 2 9B DeltaZip (2$\star$) 88.68 50.95 33.94 7.50$\times$
Table 1. Model quality of DeltaZip vs. uncompressed (FP16) vs. SparseGPT (frantar_sparsegpt_2023, ) and AWQ (AWQ, ). For Pythia, T1, T2, T3 show the accuracy on 3 downstream tasks (Amazon Review Classification, Synthetic Palindrome Numbers, Yes/No Question) from natural instructions (mishra_cross-task_2022, ). For other models, T1, T2, T3 show accuracy on 3 standard benchmarks (BoolQA, TruthfulQA, LogiQA) in lm-eval-harness (eval-harness, ). $\star$ indicates 50% structured pruning in addition to quantization.
Workload traces. To evaluate the serving performance, we use the prompts and responses sampled from the LMSys Chatbot Arena (zheng_judging_2023, ). We consider three types of model popularity distribution: 1) Uniform: all models are equally popular. 2) Skewed: model popularity follows a Zipf-$\alpha$ distribution. In other words, the popularity of the $i$-th model is proportional to $1/i^{\alpha}$. We set $\alpha=1.5$ for the skewed distribution. 3) Azure trace: since there are no publicly available traces for multi-variant LLM serving available, following previous work (li_alpaserve_2023, ), we use the Azure serverless function traces (shahrad_serverless_2020, ; zhang_faster_2021, ) as a proxy. Note the traffic in this trace is highly bursty and the model distribution is highly skewed. Except for the small-scale timeline analysis in Figure 16 (which we show for illustration purposes) and ablation studies, we run the traces for 5 minutes under different arrival rates and model distribution. Unless otherwise stated, there are 32 model variants that need to be served. The requests are sent to the serving system with a varying Poisson arrival rate (applied to the entire system).
Metrics. We use downstream accuracy to evaluate the post-compression quality and three metrics for serving performance: 1) end-to-end (E2E) latency averaged over all requests, 2) time to first token (TTFT) averaged over all requests, which is an important indicator of the system’s responsiveness, 3) throughput and 4) SLO attainment (i.e., percentage of requests served within a given SLO requirement) for TTFT and E2E Latency.
@@
System Setup. We evaluate our serving system without tuning the maximal number of concurrent deltas, since it requires offline profiling and apriori knowledge of the workload. Instead, we show two representative configurations to reveal the system performance under different settings without explicit tuning.